OpenStudy (anonymous):

The length of a rectangle is 3 times its width. The area is 56 sq cm. Find the dimensions of the rectangle.

OpenStudy (anonymous):

l=3w A=56 A=l x w, substitute l for w A=3w x w, A=3w^2 solve for w w=4.3 plug into l=3w l=13 13 x 4.3 = 55.9

OpenStudy (anonymous):

you flipped them, if length is three times width then length cannot be shorter than width

OpenStudy (anonymous):

My mistake: Area=width*length 56=3x*x 56=3x^2 18.67=x^2 x=4.32 therefore: length=3*4.32 length=12.96cm and width= 4.32cm